The length of a rectangle is twice the width. If the area is 288 square inches, what are the length and width?
Answer by mananth(16946)

width=x
length =2x
Area = l*w
=x*2x
=2x^2
2x^2=288
/2
x^2=144
x= +/-12
x=12 ignore negative
widthe =12
length =2x=24